OpenStudy (jenniferjuice):

Solve the compound inequality 30 < 10u < 120.

OpenStudy (jenniferjuice):

3 < u < 4 3 > u > 12 3 > u > 4 3 < u < 12

OpenStudy (anonymous):

Divide through by 10, since all terms are multiples of 10. 30 < 10u < 120. 3 < u < 12 Does that make sense?

OpenStudy (anonymous):

when solving these . what you do in the middle you do to the outside so i.e \[\frac{ 30 }{ 10 } < \frac{ 10 }{ 10 }u<\frac{ 120 }{10 }\] then you will have your answer :)